THE SMOOTHNESS OF RIEMANNIAN SUBMERSIONS WITH NONNEGATIVE SECTIONAL CURVATURE

JIANGUO CAO MEI-CHI SHAW

Abstract. In this article, we study the smoothness of Riemannian submersions for open manifolds with non-negative sectional curvature.